# Learning New Tasks from a Few Examples with Soft-Label Prototypes

### v3

@@ 6 Baselines
The model is optimised using negative log-likelihood. We use BERT as the encoder and fine-tune it (all layers) using our few-shot training sets and a learning rate of 3e-5.
7 Experimental setup
@@
For DeepSLP, we find that more epochs are needed to train models with a higher number of soft labels (i.e., a higher number of classes in the output of the classifier head). In other words, 3 classes fitted on a line need more epochs compared to 2 classes fitted on a line. We use AdamW Loshchilov and Hutter (2017) as our optimiser and perform hyperparameter tuning on the validation set. We find we only need a few epochs to generalise well – these range from 5 to 10 depending on the training task. We evaluate our approach and all baselines in the following manner: we fix a random seed, train our models, and evaluate performance on the test tasks. We repeat this process across three different seeds, and report the mean and standard deviation at test time.
8 Results and discussion
Category (Classes) Shot 1NN ProtoNet BERTfine-tuned* SLPBERT DeepSLPBERT
Rating DVD (3) 4 38.289 ± 6.846 37.665 ± 7.184 32.22 ± 08.72 32.298 ± 16.263 39.566 ± 5.086
8 37.598 ± 5.364 37.008 ± 5.118 36.35 ± 12.50 32.644 ± 16.016 38.788 ± 4.449
16 37.461 ± 4.506 39.123 ± 6.004 42.79 ± 10.18 35.587 ± 17.445 40.53 ± 4.375
Rating Electronics (3) 4 37.001 ± 5.164 33.696 ± 5.55 39.27 ± 10.15 35.188 ± 16.211 39.977 ± 5.959
8 37.646 ± 4.223 37.297 ± 5.938 28.74 ± 08.22 29.624 ± 12.876 41.926 ± 3.985
16 38.633 ± 3.103 43.825 ± 5.946 45.48 ± 06.13 29.836 ± 12.753 44.917 ± 3.164
Political Bias (2) 4 52.245 ± 4.348 51.15 ± 2.454 54.57 ± 5.02 53.447 ± 3.281 53.251 ± 4.042
8 54.568 ± 3.015 56.568 ± 4.228 56.15 ± 3.75 55.824 ± 3.725 58.209 ± 5.198
16 55.884 ± 2.436 59.183 ± 4.706 60.96 ± 4.25 58.277 ± 4.128 61.479 ± 2.974
Rating Kitchen (3) 4 36.769 ± 7.738 35.914 ± 6.678 34.76 ± 11.20 28.253 ± 15.907 39.624 ± 6.787
8 37.435 ± 6.347 38.46 ± 11.124 34.49 ± 08.72 24.397 ± 11.961 41.081 ± 6.777
16 38.047 ± 3.226 46.546 ± 8.394 47.94 ± 08.28 31.926 ± 18.29 45.801 ± 4.562
Political Audience (2) 4 51.827 ± 1.754 50.976 ± 1.84 51.02 ± 1.72 51.305 ± 2.68 51.741 ± 2.827
8 53.113 ± 2.314 52.022 ± 3.964 52.80 ± 2.72 53.104 ± 3.669 54.506 ± 3.274
16 53.287 ± 1.874 54.024 ± 3.071 58.45 ± 4.98 53.888 ± 3.305 56.956 ± 3.045
Sentiment Kitchen (2) 4 60.17 ± 3.197 55.24 ± 3.427 56.93 ± 7.10 61.96 ± 4.594 60.76 ± 4.426
8 59.82 ± 2.703 62.28 ± 5.103 57.13 ± 6.60 64.83 ± 3.983 65.733 ± 3.198
16 61.85 ± 2.65 66.9 ± 5.441 68.88 ± 3.39 68.21 ± 3.298 69.18 ± 2.589
Disaster (2) 4 53.629 ± 9.391 51.474 ± 8.848 55.73 ± 10.29 52.77 ± 10.803 54.252 ± 9.843
8 57.625 ± 7.638 60.661 ± 4.991 56.31 ± 09.57 56.888 ± 11.139 61.3 ± 7.961
16 60.93 ± 4.997 63.893 ± 6.62 64.52 ± 08.93 65.907 ± 3.691 69.28 ± 2.358
Airline (3) 4 46.175 ± 4.922 44.167 ± 10.752 42.76 ± 13.50 36.243 ± 22.607 50.987 ± 4.936
8 48.039 ± 5.677 50.148 ± 13.429 38.00 ± 17.06 44.972 ± 22.584 55.209 ± 6.049
16 51.754 ± 2.954 54.8 ± 10.49 58.01 ± 08.23 29.238 ± 17.494 60.247 ± 4.577
Emotion (13) 4 8.474 ± 1.028 8.953 ± 2.052 09.20 ± 3.22 8.662 ± 6.213 9.706 ± 1.108
8 8.418 ± 0.792 10.857 ± 3.436 08.21 ± 2.12 8.16 ± 3.266 8.041 ± 2.797
16 8.758 ± 0.854 11.479 ± 2.96 13.43 ± 2.51 8.115 ± 3.66 10.919 ± 1.615
Rating Books (3) 4 39.324 ± 4.985 37.715 ± 5.801 39.42 ± 07.22 25.562 ± 15.207 42.116 ± 4.725
8 38.447 ± 4.07 38.518 ± 5.327 39.55 ± 10.01 34.026 ± 14.123 42.156 ± 4.608
16 41.107 ± 2.683 44.694 ± 7.797 43.08 ± 11.78 32.509 ± 16.132 46.513 ± 3.036
Political Message (9) 4 15.044 ± 1.556 13.888 ± 2.076 15.64 ± 2.73 12.438 ± 1.799 14.421 ± 1.095
8 15.666 ± 1.053 16.155 ± 2.316 13.38 ± 1.74 15.08 ± 2.925 16.919 ± 1.756
16 14.844 ± 1.426 18.324 ± 2.011 20.67 ± 3.89 13.121 ± 3.294 18.319 ± 1.74
Scitail (2) 4 53.66 ± 4.594 50.227 ± 5.69 58.53 ± 09.74 52.296 ± 4.366 54.101 ± 3.759
8 53.212 ± 3.028 54.196 ± 6.678 57.93 ± 10.70 55.964 ± 5.705 56.341 ± 5.786
16 54.53 ± 4.399 57.744 ± 5.696 65.66 ± 06.82 59.675 ± 4.033 59.692 ± 4.227
Sentiment DVD (2) 4 53.78 ± 1.316 51.06 ± 3.302 54.98 ± 3.96 56.06 ± 2.408 55.003 ± 2.936
8 53.8 ± 2.843 55.19 ± 3.298 55.63 ± 4.34 56.98 ± 3.299 57.527 ± 3.562
16 54.05 ± 1.659 59.45 ± 3.84 58.69 ± 6.08 58.95 ± 2.813 60.76 ± 2.944
Sentiment Books (2) 4 57.56 ± 4.571 55.53 ± 4.097 54.81 ± 3.75 59.89 ± 5.385 58.67 ± 4.753
8 60.76 ± 4.177 58.97 ± 4.909 53.54 ± 5.17 64.34 ± 2.565 64.78 ± 2.615
16 60.98 ± 2.857 65.5 ± 7.026 65.56 ± 4.12 66.36 ± 2.183 67.453 ± 3.085
Restaurant (8) 4 48.194 ± 4.881 18.161 ± 2.822 49.37 ± 4.28 24.932 ± 17.102 47.634 ± 5.237
8 57.36 ± 3.681 32.146 ± 5.785 49.38 ± 7.76 29.787 ± 9.573 55.912 ± 4.494
16 64.706 ± 2.913 40.435 ± 3.348 69.24 ± 3.68 29.154 ± 13.537 61.716 ± 2.208
CoNLL (4) 4 45.104 ± 7.724 35.438 ± 7.324 50.44 ± 08.57 27.02 ± 7.346 52.724 ± 5.84
8 46.897 ± 4.183 44.259 ± 4.886 50.06 ± 11.30 31.296 ± 17.487 60.374 ± 3.731
16 53.969 ± 3.252 52.116 ± 5.354 74.47 ± 03.10 22.923 ± 7.933 67.496 ± 4.551
Table 1: Classification performance (accuracy) of our methods (SLP and DeepSLP) and baselines. Entries in green indicate the best model; * is the baseline as reported in Bansal et al. (2019). Subscripts for SLP and DeepSLP refer to the (non-fine-tuned) encoder used.

### v4

@@ 5       Partition $T_{i}$ into $D^{s}_{i}$ and $D^{q}_{i}$
Model optimisation is done using the loss function $J(\mathbf{\theta})=-log(p(y=c^{*}|x,\theta))$.
A.3 Analysis of soft-labels derived from linear constraints: constraintSLP
Theorem A.1.
The soft-label value of each class within a single soft-label prototype generated using constraintSLP is inversely proportional to its distance from the soft-label prototype along the line connecting all classes captured by it.
Proof of Theorem A.1 (Informal)
Consider a line $l$ connecting three class centroids (while we focus on a three class system, the conclusions generalise to $n>3$ classes too). The class centroids are represented by A, B and C. The soft-label prototypes at ends A and C contain the values $[a_{1},a_{2},a_{3}]$ and $[c_{1},c_{2},c_{3}]$ respectively. Consider a support example $x\in A$ at a distance $d_{a}$ and $d_{b}$ from A and C respectively. Directly using the constraints in Algorithm 4 of Sucholutsky et al. (2021), we state that the influence of A (i.e., the distance-weighted sum of the soft-labels at $x$) should be more than the sum of the influence of the other two classes. Thus, we need to maximise:
$\frac{a_{1}}{d_{a}}+\frac{c_{1}}{d_{b}}>\left(\frac{a_{2}}{d_{a}}+\frac{c_{2}}{d_{b}}\right)+\left(\frac{a_{3}}{d_{a}}+\frac{c_{3}}{d_{b}}\right)$ (2)
As we move $x$ further towards A, $d_{a}\rightarrow 0$ and the influence of $[a_{1}/d_{1},a_{2}/d_{2},a_{3}/d_{3}]$ increases thus $\sum^{i=3}_{i=2}c_{i}/d_{b}<<<\sum^{i=3}_{i=2}a_{i}/d_{a}$. Therefore we have the approximation:
$\centering\frac{a_{1}}{d_{a}}>\frac{a_{2}}{d_{a}}+\frac{a_{3}}{d_{a}}\implies a_{1}>a_{2}+a_{3}\@add@centering$ (3)
If we take a support example $x\in C$, by symmetry as $x$ is moved towards C, $d_{b}\rightarrow 0$, we can also write:
$\centering\frac{c_{3}}{d_{b}}>\frac{c_{2}}{d_{b}}+\frac{c_{1}}{d_{b}}\implies c_{3}>c_{2}+c_{1}\@add@centering$ (4)
Furthermore, consider a point $x$ in the middle of $l$ equidistant from A and C (by a distance $d$) – such a point will always be classified as $x\in B$. Thus, the influence of B should be higher than both A and C. Thus we have:
$\frac{a_{2}}{d}+\frac{c_{2}}{d}>\frac{a_{1}}{d}+\frac{c_{1}}{d}\;\;\;\&\;\;\;\frac{a_{2}}{d}+\frac{c_{2}}{d}>\frac{a_{3}}{d}+\frac{c_{3}}{d}$
$\implies a_{2}+c_{2}>a_{1}+c_{1}\;\;\;\&\;\;\;a_{2}+c_{2}>a_{3}+c_{3}$
From Equation 3 and Equation 4 we can replace $a_{1}$ and $c_{3}$ and get:
$a_{2}+c_{2}>a_{2}+a_{3}+c_{1}\implies c_{2}>c_{1}$
$a_{2}+c_{2}>a_{3}+c_{2}+c_{1}\implies a_{2}>a_{3}$
Therefore, we have:
$a_{1}>a_{2}>a_{3}\;\;\;\&\;\;\;c_{3}>c_{2}>c_{1}$
This is an intuitive result as the soft-label value of each class decreases as the distance of the class centroid increases from the prototype location – the class nearest to the prototype has the highest soft-label value and the class furthest away has the lowest soft-label value.
Recall that $\sum_{i=1}^{3}a_{i}=1$ and $\sum_{i=1}^{3}c_{i}=1$ and $a_{i},c_{i}\geq 0$ $\forall i=\{1,2,3\}$ otherwise the optimisation problem becomes unbounded. Therefore, the ranges of values for $[a_{1},a_{2},a_{3}]$ and $[c_{1},c_{2},c_{3}]$ are:
$a_{3}\in[0,a_{2}),\;a_{2}\in(a_{3},a_{1}),\;a_{1}\in(a_{2},1]$
$c_{3}\in(c_{2},1],\;c_{2}\in(c_{1},c_{3}),\;c_{1}\in[0,c_{2})$
Using Algorithm 4 in Sucholutsky et al. (2021), we see that there are multiple constraints for $a_{1}$ and $a_{2}$ which require them to be maximised, but there are none for $a_{3}$. Thus, to maximise Equation 3, $a_{3}$ adjusts to the minimum value it can get:
$a_{3}=min(0,a_{2})=\epsilon\simeq 0$
By symmetry, we can also conclude that:
$c_{1}=min(0,c_{2})=\epsilon\simeq 0$
These approximations are also generalisable to multiple classes connected by a line, for example, if a line connects only two centroids, the soft-labels at each end are derived as $[0,1]$ and $[1,0]$ the same as a “hard" label. These findings are substantiated experimentally in Table 2 where we examine the soft-labels generated by DeepSLP and constraintSLP using a few-shot training support set of the task airline with $8$ examples per class – the constraintSLP soft label corresponding to the furthest class from the prototype location drops to almost zero compared to other soft label values. On the other hand, DeepSLP prevents overfitting on the nearest classes and produces a more generalised distribution of soft-labels. This is a trend generally observed in other tasks and classes as well. We further use this theorem to prove the main theorem given by Theorem A.2.
# constraintSLP DeepSLP($x$)
1 $015.6422e-$ 01 $019.7887e-$ 01
$014.3577e-$ 01 $022.0995e-$ 02
$169.7973e-$ 16 $041.3500e-$ 04
2 $135.3090e-$ 13 $022.5240e-$ 02
$014.3212e-$ 01 $019.7475e-$ 01
$015.6787e-$ 01 $051.0000e-$ 05
Table 2: Soft labels derived using constraintSLP and DeepSLP. # denotes the index of the soft-label prototype lying on the line. Soft labels are constant for constraintSLP, however, they are a function of input point $x$ for DeepSLP, thus allowing more flexibility.
Theorem A.2.
The constant soft-labels in constraintSLP do not always select the closest class centroid to a test point.
Proof of Theorem A.2 (Informal)
Furthermore, consider a line $b$ perpendicular to $l$ – it intersects $l$ between A and B. We select $\theta$ such that $\phi=\pi/2$. We denote the complete setup diagrammatically in Figure 5.
Figure 5: Schematic diagram for ascertaining $\theta$ with class centroids $A=(-10,0)$, $B=(5,0)$ and $C=(15,0)$.
Consider the distance weighted influences at $x$ for class A. We have the influence as $\frac{a_{1}sin\theta}{e}+\frac{c_{1}cos\theta}{e}=\frac{a_{1}sin\theta}{e}$ as $c_{1}\simeq 0$. Similarly, for class B, we have the weighted influence as $\frac{a_{2}sin\theta}{e}+\frac{c_{2}cos\theta}{e}$. To calculate values of $\theta$ where the weighted influence of B is more than the weighted influence of A, we get:
$\frac{a_{2}sin\theta}{e}+\frac{c_{2}cos\theta}{e}>\frac{a_{1}sin\theta}{e}$
$\implies a_{2}sin\theta+c_{2}cos\theta>a_{1}sin\theta$
$\implies c_{2}cos\theta-(a_{1}-a_{2})sin\theta>0$
$\implies\frac{c_{2}cos\theta-(a_{1}-a_{2})sin\theta}{\sqrt{c_{2}^{2}+(a_{1}-a_{2})^{2}}}>0$
$\implies cos(\theta+\alpha)>0$
(a) Denoting the decision boundaries calculated with constraintSLP. Green represents points classified as class A, pink represents the points classified as class B, and brown represents the points classified as C.
(b) Zooming in at point $x$. We can see that it is classified as B.
Figure 6: The soft-labels of the linear constraint system at A and C using constraintSLP are calculated as $[0.5963,0.4036,0.0001]$ and $[0.0001,0.4495,0.5504]$. We also get $\theta=66.84^{\circ}$. Using these soft-labels, we calculate the decision boundaries for points in this area. We use $\theta$ to calculate the coordinates of $x$. Zooming in, we can visually inspect that $x$ is classified to class B. For $x$, the Euclidean distance of $x$ from A and B is $9.847$ and $14.338$ respectively. From the figure, we can see that constraintSLP classifies $x$ as B even though the Euclidean distance of $x$ from A is shorter.
where $\alpha=tan^{-1}\left(\frac{a_{1}-a_{2}}{c_{2}}\right)$. Since $cos(\theta+\alpha)>0$ we have $(\theta+\alpha)\in(-\pi/2,\pi/2)$ and since $\theta>0$, thus for $\theta\in[0,\pi/2-\alpha)$, the weighted influence of B is more than the weighted influence of A.
However, it is worth observing the result for $\theta$ derived above can contain points closer to A (using Euclidean distance) which are actually classified as B. We can easily demonstrate this with a counter example explained in Figure 6.
Therefore, for points closer to A compared to B using an Euclidean measure, constraintSLP can still return a higher value for the influence at B compared to A. This adversely affects performance in classifiers where we rely on selection of the closest class centroid for classification – such as 1-NN, Prototypical Networks, and constraintSLP – and we believe this is the reason behind the poor performance of constraintSLP for cases where the total classes is greater than two.
A.4 Training algorithms for MetaSLP
Inner-loop training
Our inner-loop training algorithm for MetaSLP is presented in Algorithm 3. The inner-loop encoder optimisation can be understood as updating the parameters of the encoder to “push” different classes away from each other and “pull” points belonging to the same class together; i.e., increase inter-class distance and decrease intra-class distance which leads to well-defined clusters per class. We present this process in Figure 7.
Figure 7: Inner-loop training – note that inter-class embeddings are pushed further away, and intra-class embeddings are pushed closer together across $n$ steps. The endpoints of the line mark the location of the soft-label prototypes.
1 $\mathcal{T}_{i}\leftarrow$ meta-training task
2 $\mathcal{L}\leftarrow$ line connecting both centroids of a task
3 $\alpha\leftarrow$ inner-loop learning rate
4 $\mathcal{S}\leftarrow$ inner-loop optimisation steps
5 Initialise $g_{1}x(x)$ and $g_{2}(x)$ randomly
6 while $s<\mathcal{S}$ do
7       Sample support examples $X^{s}$ for $\mathcal{T}_{i}$
8       Calculate locations of each soft-label prototype in $\mathcal{L}$
9       Use Equation 1 to classify $x\in X^{s}$
10       Calculate $\nabla\mathcal{L}_{\mathcal{T}_{i}}(f_{\theta}(x),g_{1}(x),g_{2}(x))$
11       Scale $\nabla_{\phi}\mathcal{L}_{\mathcal{T}_{i}}(g_{1}(x))$ and $\nabla_{\omega}\mathcal{L}_{\mathcal{T}_{i}}(g_{2}(x))$ by the distances from the soft-label prototypes
12       $\theta^{\prime}_{i}\leftarrow\theta-\alpha\nabla_{\theta}\mathcal{L}_{\mathcal{T}_{i}}(f_{\theta}(x))$
13       $\theta^{\prime}_{1_{i}}\leftarrow\theta_{1}-\alpha\nabla_{\phi}\mathcal{L}_{\mathcal{T}_{i}}(g_{1}(x))$
14       $\theta^{\prime}_{2_{i}}\leftarrow\theta_{2}-\alpha\nabla_{\omega}\mathcal{L}_{\mathcal{T}_{i}}(g_{2}(x))$
15 end while
Algorithm 3 Inner-loop training of MetaSLP
Outer-loop training
Our outer-loop training algorithm is presented in Algorithm 4.
1 $\mathcal{T}\leftarrow$ batch of meta-training tasks, $|\mathcal{T}|=n$
2 $\mathcal{M}\leftarrow$ batch of distinct inner-loop optimised models parameterised by $\theta_{i}$, $|\mathcal{M}|=n$
3 $\beta\leftarrow$ outer-loop learning rate
4 for $\mathcal{T}_{i},\;\mathcal{M}_{i}\in\mathcal{T},\;\mathcal{M}$ do
5       if FOMAML then
6             Sample query examples $X^{q}$ for $\mathcal{T}_{i}$, $X^{s}\cap X^{q}=\Phi$
7             Use Equation 1 to classify $x\in X^{q}$
8             Calculate $\nabla_{\theta^{\prime}_{i}}\mathcal{L}_{\mathcal{T}_{i}}(f_{\theta_{i}^{\prime}}(x),g_{\phi_{i}^{\prime}}(x),h_{\omega_{i}^{\prime}}(x))$
9             Update $\nabla_{\theta}\mathcal{L}_{\mathcal{T}_{i}}(f_{\theta}(x))+=\nabla_{\theta^{\prime}_{i}}\mathcal{L}_{\mathcal{T}_{i}}(f_{\theta_{i}^{\prime}}(x),g_{\phi_{i}^{\prime}}(x),h_{\omega_{i}^{\prime}}(x))$
10       end if
11      if Reptile then
12             Update $\nabla_{\theta}\mathcal{L}_{\mathcal{T}_{i}}(f_{\theta}(x))+=\theta-\theta_{i}$
13       end if
14
15 end for
Update $\theta\leftarrow\theta-\frac{\beta}{n}\sum^{n}_{i=1}\nabla_{\theta}\mathcal{L}_{\mathcal{T}_{i}}(f_{\theta}(x))$
Algorithm 4 Outer-loop training of MetaSLP
A.5 Meta-training details
GLUE Wang et al. (2018) tasks and their details are provided in Table 3. These tasks include MNLI Williams et al. (2018), SST2 Socher et al. (2013), CoLA Warstadt et al. (2018), MRPC Dolan and Brockett (2005), QQP Wang et al. (2017), QNLI Wang et al. (2018), RTE Giampiccolo et al. (2008) and SNLI Bowman et al. (2015). We employ the same tasks as Bansal et al. (2020a) to ensure direct comparability. Note that the datasets and classes in GLUE are completely different from the datasets used for evaluating the model - thus the final model fine-tunes on unseen few-shot data and learns classes it has previously not encountered.
To train our model to detect the sentiment contained within phrases of a sentence by using the annotations for phrases within sentences for SST2, we append a separator token and the annotated phrase for each sentence at the end of the sentence in the form “[CLS] <sentence_1> [SEP] <sentence_2> [SEP]" and obtain the passage level embedding for training.
Dataset Labels Training Size Validation Size Test Size
CoLA 2 8551 1042 -
MRPC 2 3669 409 -
QNLI 2 104744 5464 -
QQP 2 363847 40431 -
RTE 2 2491 278 -
SNLI 3 549368 9843 -
SST-2 2 67350 873 -
MNLI 3 392703 19649 -
Table 3: Details of GLUE tasks used for meta-training.
A.6 Hyperparameters
@@
For DeepSLP, we find that more epochs are needed to train models with a higher number of soft labels (i.e., a higher number of classes in the output of the classifier head) - essentially, 3 classes fitted on a line need more epochs compared to 2 classes fitted on a line. We use AdamW Loshchilov and Hutter (2017a) as our optimiser and perform hyperparameter tuning on the validation set. We only need a few epochs (5 to 10) to generalise well depending on the training task. We fix a random seed, train our models and evaluate performance on the test tasks. We repeat this process across three different seeds and report the mean and standard deviation. Hyperparameters for all baselines in this setting can be found online in our code repository [^5] https://github.com/avyavkumar/meta-learned-lines.
Meta-training
For inner-loop optimisation, we use SGD as an optimiser with Nesterov and a momentum factor. We use a cosine annealing learning rate scheduler (Loshchilov and Hutter, 2017b) on our outer-loop learning rate to decay the learning rate from a starting rate to an end rate without restarts across one epoch for MetaSLPREPTILE. We use AdamW (Loshchilov and Hutter, 2017a) as our outer-loop optimiser with AMSGrad (Reddi et al., 2018). We employ early stopping and stop training if our model does not improve it’s validation set accuracy over 100 batches. We use learnable inner-loop learning rates for MetaSLPFOMAML per parameter group for better optimisation as indicated by previous literature Antoniou et al. (2019). All meta-training hyperparameters can be found in Appendix A.6, Table 4.
Meta-testing
Similar to inner-loop optimisation at meta-training, we use SGD with Nesterov and the same optimiser hyperparameters. However, we decay the learning rate using a cosine scheduler across all fine-tuning epochs to prevent overfitting on the few-shot (support) training set per task for MetaSLPREPTILE.
Parameter Search Space MetaSLPREPTILE MetaSLPFOMAML Reptile
Tunable layers ($v$) $[1,2,3,4]$ $4$ $4$ $4$
K-shot $[8,16,32]$ $16$ $16$ $16$
Batch size $[8,16,32]$ $16$ $16$ $16$
Steps $[3,5,7]$ $5$ $5$ $5$
$\alpha_{f}$ $[5e-3,1e-3,1e-4]$ $35e-$ 3 Learnable $31e-$ 3
$\alpha_{g},\;\alpha_{h}$ $[5e-3,1e-3,1e-4,1e-2]$ $35e-$ 3 Learnable $31e-$ 3
Nesterov $[True,False]$ $True$ $True$ $True$
Momentum $[0.5,0.7,0.9]$ $0.9$ $0.9$ $True$
$\beta$initial $[1e-5,2e-5,5e-5]$ $55e-$ 5 $52e-$ 5 $51e-$ 5
$\beta$final $[1e-5,2e-5,5e-6]$ $52e-$ 5 $52e-$ 5 $51e-$ 5
Task sampling [square root, uniform] square root square root square root
Table 4: Meta-training hyperparameters.
A.7 Results
The complete set of results of all models and baselines can be seen in Table 5 for the low-resource setting and DeepSLP, and Table 6 for the high-resource setting and MetaSLP.
DeepSLP